### v1

@@ 2.4 Conjectured properties of the typing judgment
  1. 1.
    (IsDefEqU.trans) If $\Gamma\vdash e_{1}\equiv e_{2}$ and $\Gamma\vdash e_{2}\equiv e_{3}$, then $\Gamma\vdash e_{1}\equiv e_{3}$.
    (This is t-trans but without requiring the types to match.)
  2. 2.
    (isDefEq_iff) $\Gamma\vdash e\equiv e^{\prime}:\alpha$ if and only if $\Gamma\vdash e:\alpha$, $\Gamma\vdash e^{\prime}:\alpha$, and $\Gamma\vdash e\equiv e^{\prime}$.
    (Hence this formulation implies the one of [4].)
  3. 3.
    (IsDefEqU.defeqDF) If $\Gamma\vdash e\equiv e^{\prime}:\alpha$ and $\Gamma\vdash\alpha\equiv\beta$, then $\Gamma\vdash e\equiv e^{\prime}:\beta$.
    (This is t-conv but with an untyped definitional equality.)
This group of theorems is proved mutually with 2.13 in [4]:
@@
  2. 2.
    (IsDefEqU.forallE_inv)
    If $\Gamma\vdash(\forall x:\alpha.\;\beta)\equiv(\forall x:\alpha^{\prime}.\;\beta^{\prime})$ then $\Gamma\vdash\alpha\equiv\alpha^{\prime}$ and $\Gamma,x:\alpha\vdash\beta\equiv\beta^{\prime}$.
  3. 3.
@@
The reason 2.13 and 2.15 have been downgraded from theorems in [4] to conjectures here is because the proof has an error in one of the technical lemmas, and it remains to be seen if it is possible to salvage the proof. More precisely, the proof constructs a stratification $\vdash_{i}$ of the typing judgment in order to break the mutual induction between typing and definitional equality, but this stratification does not and cannot respect substitution; that is, if $\Gamma,x:\beta\vdash_{i}e:\alpha$ and $\Gamma\vdash_{j}e^{\prime}:\beta$, then the proof requires $\Gamma\vdash_{\max(i,j)}e[x\mapsto e^{\prime}]:\alpha$ but only $\Gamma\vdash_{i+j}e[x\mapsto e^{\prime}]:\alpha$ holds.
We still have reasons to believe the conjectures here are true, but more work is needed to determine how to structure the proof by induction. The proof of soundness is not impacted because there are alternative routes to construct the model that avoid unique typing (also described in [4]), but these conjectures are necessary in at least some form in order to prove the correctness of the typechecker (see section 3).
@@
Conjecture 2.16 (Reverse weakening).
If $e,e^{\prime}$ do not mention any variables in $\Delta$, then
$\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}$ if and only if $\Gamma,\Gamma^{\prime}\vdash e\equiv e^{\prime}$.
Corollary 2.17 (Consequences of reverse weakening).
  1. 1.
    If $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$ and $e,e^{\prime}$ do not mention the variables in $\Delta$, then there exists $\alpha^{\prime}$ not mentioning $\Delta$ such that $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha^{\prime}$.
  2. 2.
    If $e,e^{\prime},\alpha$ do not mention any variables in $\Delta$, then $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$ if and only if $\Gamma,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$.
  3. 3.
    If $\vdash\Gamma,\Delta,\Gamma^{\prime}\ \mathsf{ok}$ and $\Gamma^{\prime}$ does not mention the variables in $\Delta$ then $\vdash\Gamma,\Gamma^{\prime}\ \mathsf{ok}$.
@@
3 The typechecker
3.1 Relating VExpr to Expr
@@
  | bvar (deBruijnIndex : Nat)
  | fvar (fvarId : FVarId)
  | mvar (mvarId : MVarId)
  | sort (u : Level)
  | const (declName : Name) (us : List Level)
  | app (fn arg : Expr)
  | lam (binderName : Name) (binderType body : Expr) (binderInfo : BinderInfo)
  | forallE (binderName : Name) (binderType body : Expr) (binderInfo : BinderInfo)
  | letE (declName : Name) (type value body : Expr) (nonDep : Bool)
  | lit : Literal  Expr
  | mdata (data : MData) (expr : Expr)
  | proj (typeName : Name) (idx : Nat) (struct : Expr)
@@
  • The Level type (not shown) used in the sort constructor also differs in similar ways to VLevel, containing a mvar constructor we don’t care about and a param constructor which takes a Name instead of a Nat.
@@
  • proj \$c\$ \$i\$ \$e\$ is a projection out of a structure or structure-like inductive type. (See section 5.) This is the most challenging constructor to desugar, because even though it is equivalent to an application of the recursor for the inductive type, we need the type of $e$ to construct this term. This means that the desugaring function must be type-aware, and is only well-defined up to definitional equality.
@@
inductive TrExpr (env : VEnv) (Us : List Name) : VLCtx  Expr  VExpr  Prop
@@
  | vlam (type : VExpr)          x : type
  | vlet (type value : VExpr)    x : type := value
@@
  | (_, vlam α) :: Δ => α :: toCtx Δ
  | (_, vlet _ _) :: Δ => toCtx Δ
@@
The Option FVarId in the context represents the “name” of the variables. Lean follows the “locally nameless” discipline, in which most operations act only on “closed terms” (where closed means that there are no bvar \$i\$ variables outside a binder), so any time it needs to process a term under a binder it first instantiates the term, replacing bvar \$(i+1)\$ with bvar \$i\$ and bvar 0 with fvar \$a\$, where \$a\$ : FVarId is a (globally) fresh variable name. We store these names in $\Delta$ so that we know how to relate them to de Bruijn indices in the context.
@@ 3.2 The typechecker implementation
def inferLambda (e : Expr) (inferOnly : Bool) : RecM Expr := loop #[] e where
@@
    withLCtx ((← getLCtx).mkLocalDecl id name d bi) do
@@
        _  ensureSortCore (← inferType d inferOnly) d
@@
  | e => do
    let r  inferType (e.instantiateRev fvars) inferOnly
@@
expr type_checker::infer_lambda(expr const & _e, bool infer_only) {
@@
    expr d = instantiate_rev(binding_domain(e), fvars.size(), fvars.data());
    expr fvar = m_lctx.mk_local_decl(m_st->m_ngen, binding_name(e), d, binding_info(e));
@@
      ensure_sort_core(infer_type_core(d, infer_only), d);
@@
  expr r = infer_type_core(instantiate_rev(e, fvars.size(), fvars.data()), infer_only);
@@
Although control flow is expressed differently (here, the replacement of a while loop with a tail-recursive function), the two implementations do essentially the same things in the same order. As a result, most bugs or peculiarities we have discovered in the course of formalizing Lean4Lean are replicated in the C++ kernel, which gives us hope that verifying Lean4Lean will lend strong evidence to the correctness of the C++ code as well.
3.2.1 Termination and the RecM monad
@@
  whnfCore (e : Expr) (cheapRec := false) (cheapProj := false) : M Expr
@@
  • Mutual definitions generally compile to a well founded recursion, while we want to utilize structural recursion when possible because of the better induction principles.
@@
  • As we will see, there is no suitable well founded measure in the first place, so we would be forced to declare the definition as partial, but we absolutely cannot do this if we want any hope of proving correctness because it would make our functions opaque, so we need another option.
@@
  • whnfCore \$e\$ cheapRec cheapProj has the same specification as whnf \$e\$. cheapRec and cheapProj are flags affecting which kinds of terms are unfolded. The reason both whnf and whnfCore show up separately in this list is because they are both used in other functions.
  • inferType \$e\$ inferOnly infers or type-checks a term. That is, if inferType \$e\$ inferOnly returns $\alpha$ then $\Gamma\vdash e:\alpha$ holds if either $e$ is well typed or inferOnly is false.
@@
  • It is unlikely that we can prove termination of a typechecker for Lean in Lean, because although our soundness proof does not depend on termination, MetaCoq’s does [11], and generally termination measures for DTT require large cardinals of comparable strength to the proof theory. We are up against Gödel’s incompleteness theorem, so anything that would imply the unconditional soundness of Lean won’t be directly provable.
  • Besides this, the Lean type theory is known not to terminate. Coquand and Abel [1] constructed a counterexample to strong normalization using reduction of proofs, and this can be shown to impact definitional equality checks even for regular types:
@@
       @cast (True  True’) A (propext fun _ => a, fun _ => id⟩) <|
       fun z => z (True  True’) id z
@@
    example : foo.recOn (fun _ => 1) = 1 := by rfl  isDefEq nontermination
@@
  • The kernel does not loop forever in many cases because this is a bad user experience — it has timeouts and depth limits. Not all parts of the kernel have such limits, but it does give us a reasonable design principle which fortuitously solves our termination problem.
@@ 4 The global environment
  • add : Environment  ConstantInfo  Environment – add a ConstantInfo declaration to the environment
  • mkEmptyEnvironment : IO Environment – constructs an environment with no constants [^6] This is in IO because environment extensions can perform computation in IO for their initialization. But the constant map, which is what we care to verify, is initialized to the empty map.
  • find? : Environment  Name  Option ConstantInfo – retrieves a declaration by name
@@
structure VConstant where (uvars : Nat) (type : VExpr)
structure VDefEq where (uvars : Nat) (lhs rhs type : VExpr)
@@
5 Inductive types
Lean’s implementation of inductive types is based on Dybjer [8]. It notably differs from Coq in that rather than having primitive fix and match expressions, we have constants called “recursors” generated from the inductive specification. This choice significantly simplifies the VExpr type, which as we have seen contains no special support for anything inductive-related except for the generic def.eq. hook t-extra. It also means we do not need a complex guard checker, something which MetaCoq currently axiomatizes ([11]) and which has been the source of soundness bugs in the past. Instead, the complexity is pushed to the generation of the recursor for the inductive type.
For single inductives, the algorithm is described in full in [4]. In Lean 3, nested and mutual inductives were simulated using single inductives so the kernel only had to deal with the single inductive case. However this simulation process was both costly and did not cover the full grammar of nested and mutual inductives, so Lean 4 moved the checking to the kernel.
For a simple example of a mutual inductive, we can define a data-carrying version of Even and Odd by mutual induction like so:
@@ 5 Inductive types
  | zero : Even 0
  | succ : Odd n  Even (n+1)
@@
@Even.rec motive_1 motive_2 F_zero F_succ F_succ 0 Even.zero  F_zero
@Even.rec motive_1 motive_2 F_zero F_succ F_succ (n+1) (Even.succ e)
  @F_succ n e (@Odd.rec motive_1 motive_2 F_zero F_succ F_succ n e)
@Odd.rec motive_1 motive_2 F_zero F_succ F_succ (n+1) (Odd.succ e)
  @F_succ n e (@Even.rec motive_1 motive_2 F_zero F_succ F_succ n e)
@@
  {motive_1 : T  Sort u}  {motive_2 : List T  Sort u}
  ((a : List T)  motive_2 a  motive_1 (T.mk a))
  motive_2 []
@@
    motive_1 head  motive_2 tail  motive_2 (head :: tail))
  (t : T)  motive_1 t
@@
Currently, Lean4Lean contains a complete implementation of inductive types, including nested and mutual inductives and all the type-checking effects of this, but not much work has been done on the theoretical side, defining what an inductive specification should generate. Ideally we would like to do it in a way which is not so intimately tailored to the nested inductive collection process (i.e. we can see the order in which subterms are visited by the order of motive_\$n\$ arguments in the above recursor) and instead allows a range of implementation strategies which can all be proved sound against a set-theoretic model.
5.1 Eta for structures
Another new kernel feature in Lean 4 is primitive projections, and eta for structures, which says that if $s$ is a structure (which is to say, a one-constructor non-recursive inductive type), then $s\equiv\mbox{mk}(s.1,\dots,s.n)$ where $s.i$ is the $i$’th projection (Expr.proj). This was already true as a propositional equality proven by induction on $s$, but it is now a definitional equality and the kernel has to unfold structures when needed to support this. This does not fundamentally change the theory; indeed in [4] a mild form of eta for pairs was introduced in order to make the reduction to W-types work, so we believe that the main results should still hold even with this extension.
6 Results

### v2

@@ 2.4. Conjectured properties of the typing judgment
  1. (1)
    (IsDefEqU.trans) If $\Gamma\vdash e_{1}\equiv e_{2}$ and $\Gamma\vdash e_{2}\equiv e_{3}$, then $\Gamma\vdash e_{1}\equiv e_{3}$. (This is t-trans but without requiring the types to match.)
  2. (2)
    (isDefEq_iff) $\Gamma\vdash e\equiv e^{\prime}:\alpha$ if and only if $\Gamma\vdash e:\alpha$, $\Gamma\vdash e^{\prime}:\alpha$, and $\Gamma\vdash e\equiv e^{\prime}$. (Hence this formulation implies the one of (Carneiro, 2019).)
  3. (3)
    (IsDefEqU.defeqDF) If $\Gamma\vdash e\equiv e^{\prime}:\alpha$ and $\Gamma\vdash\alpha\equiv\beta$, then $\Gamma\vdash e\equiv e^{\prime}:\beta$. (This is t-conv but with an untyped definitional equality.)
This group of theorems is proved mutually with Conjecture 2.7 in (Carneiro, 2019):
@@
  2. (2)
    (IsDefEqU.forallE_inv) If $\Gamma\vdash(\forall x:\alpha.\;\beta)\equiv(\forall x:\alpha^{\prime}.\;\beta^{\prime})$ then $\Gamma\vdash\alpha\equiv\alpha^{\prime}$ and $\Gamma,x:\alpha\vdash\beta\equiv\beta^{\prime}$. [^7] This theorem is sometimes called “injectivity of $\forall$” but this name is slightly misleading because $\forall$ is not injective as a function on types, which is to say if we use propositional equality instead of def.eq then the theorem fails.
  3. (3)
@@
The reason Conjecture 2.7 and Conjecture 2.9 have been downgraded from theorems in (Carneiro, 2019) to conjectures here is because the proof has an error in one of the technical lemmas, and it remains to be seen if it is possible to salvage the proof. More precisely, the proof constructs a stratification $\vdash_{i}$ of the typing judgment in order to break the mutual induction between typing and definitional equality, but this stratification does not and cannot respect substitution; that is, if $\Gamma,x:\beta\vdash_{i}e:\alpha$ and $\Gamma\vdash_{j}e^{\prime}:\beta$, then the proof requires $\Gamma\vdash_{\max(i,j)}e[x\mapsto e^{\prime}]:\alpha$ but only $\Gamma\vdash_{i+j}e[x\mapsto e^{\prime}]:\alpha$ holds.
We still have reasons to believe the conjectures here are true, but more work is needed to determine how to structure the proof by induction. The proof of soundness is not impacted because there are alternative routes to construct the model that avoid unique typing (also described in (Carneiro, 2019)), but these conjectures are necessary in at least some form in order to prove the correctness of the typechecker (see section 3).
@@
Conjecture 2.10 (Strengthening).
If $e,e^{\prime}$ do not mention any variables in $\Delta$, then $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}$ if and only if $\Gamma,\Gamma^{\prime}\vdash e\equiv e^{\prime}$.
Corollary 2.11 (Consequences of strengthening).
  1. (1)
    (IsDefEq.skips) If $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$ and $e,e^{\prime}$ do not mention the variables in $\Delta$, then there exists $\alpha^{\prime}$ not mentioning $\Delta$ such that $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha^{\prime}$.
  2. (2)
    (IsDefEq.weakN_iff) If $e,e^{\prime},\alpha$ do not mention any variables in $\Delta$, then $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$ if and only if $\Gamma,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$.
  3. (3)
    (OnCtx.weakN_inv) If $\vdash\Gamma,\Delta,\Gamma^{\prime}\ \mathsf{ok}$ and $\Gamma^{\prime}$ does not mention the variables in $\Delta$ then $\vdash\Gamma,\Gamma^{\prime}\ \mathsf{ok}$.
@@
3. The typechecker
3.1. Relating VExpr to Expr
@@
| bvar (deBruijnIndex : Nat)
| fvar (fvarId : FVarId)
| mvar (mvarId : MVarId)
| sort (u : Level)
| const (declName : Name) (us : List Level)
| app (fn arg : Expr)
| lam (binderName : Name) (type body : Expr)
(info : BinderInfo)
| forallE (binderName : Name) (type body : Expr)
(info : BinderInfo)
| letE (declName : Name) (type value body : Expr)
(nonDep : Bool)
| lit : Literal Expr
| mdata (data : MData) (expr : Expr)
| proj (typeName : Name) (idx : Nat)
(struct : Expr)
@@
  • The Level type (not shown) used in the sort constructor also differs in similar ways to VLevel, containing an mvar constructor we don’t care about and a param constructor which takes a Name instead of a Nat.
@@
  • proj \$c\$ \$i\$ \$e\$ is a projection out of a structure or
    structure-like inductive type. (See section 5.) This is the most challenging constructor to desugar, because even though it is equivalent to an application of the recursor for the inductive type, we need the type of $e$ to construct this term. This means that the desugaring function must be type-aware, and is only well-defined up to definitional equality.
@@
inductive TrExpr (env : VEnv) (Us : List Name) :
VLCtx Expr VExpr Prop
@@
| vlam (type : VExpr) x : type
| vlet (type value : VExpr) x : type := value
@@
| (_, vlam α) :: Δ => α :: toCtx Δ
| (_, vlet _ _) :: Δ => toCtx Δ
@@
The Option FVarId in the context represents the “name” of the variables. Lean follows the “locally nameless” discipline, in which most operations act only on “closed terms” (where closed means that there are no bvar \$i\$ variables outside a binder), so any time it needs to process a term under a binder it first instantiates the term, replacing bvar \$(i+1)\$ with bvar \$i\$ and bvar 0 with fvar \$a\$, where \$a\$ : FVarId is a (globally) fresh variable name. We store these names in $\Delta$ so that we know how to relate them to de Bruijn indices in the context. The name generator uses a counter in the global state of the typechecker, and there is an invariant that all FVarIds in the state are less than the counter value.
@@ 3.2. The typechecker implementation
def inferLambda (e : Expr) (inferOnly : Bool) :
RecM Expr := loop #[] e where
@@
let lc := (← getLCtx).mkLocalDecl id name d bi
withLCtx lc do
@@
_ ensureSortCore
(← inferType d inferOnly) d
@@
| e => do
let r inferType
(e.instantiateRev fvars) inferOnly
@@
expr type_checker::infer_lambda(expr const & _e,
bool infer_only) {
@@
expr d = instantiate_rev(binding_domain(e),
fvars.size(), fvars.data());
expr fvar = m_lctx.mk_local_decl(m_st->m_ngen,
binding_name(e), d, binding_info(e));
@@
ensure_sort_core(
infer_type_core(d, infer_only), d); }
@@
expr r = infer_type_core(instantiate_rev(
e, fvars.size(), fvars.data()), infer_only);
@@
Although control flow is expressed differently (here, the replacement of a while loop with a tail-recursive function), the two implementations do essentially the same things in the same order. As a result, most bugs or peculiarities we have discovered in the course of formalizing Lean4Lean are replicated in the C++ kernel, which gives us hope that verifying Lean4Lean will lend strong evidence to the correctness of the C++ code as well. We have taken care to write things in a style which will lend itself well to the verification task.
3.2.1. Termination and the RecM monad
@@
whnfCore (e : Expr) (cheapRec := false)
(cheapProj := false) : M Expr
@@
  • Mutual definitions generally compile to a well-founded recursion, while we want to utilize structural recursion when possible because of the better induction principles.
@@
  • As we will see, there is no suitable well-founded measure in the first place, so we would be forced to declare the definition as partial, but we absolutely cannot do this if we want any hope of proving correctness because it would make our functions opaque, so we need another option.
@@
  • whnfCore \$e\$ cheapRec cheapProj has the same specification as whnf \$e\$, but it lacks some of the early exit paths. cheapRec and cheapProj are flags affecting which kinds of terms are unfolded. The reason both whnf and whnfCore show up separately in this list is because they are both used in other functions (i.e. some functions call whnfCore directly because e.g. the fast path isn’t applicable).
  • inferType \$e\$ inferOnly infers or typechecks a term. That is, if inferType \$e\$ inferOnly returns $\alpha$ then $\Gamma\vdash e:\alpha$ holds if either $e$ is well typed or inferOnly is false.
@@
  • It is unlikely that we can prove termination of a typechecker for Lean in Lean, because although our soundness proof does not depend on termination, MetaCoq’s does (Sozeau et al., 2019), and generally termination measures for DTT require large cardinals of comparable strength to the proof theory. We are up against Gödel’s incompleteness theorem, so anything that would imply the unconditional soundness of Lean won’t be directly provable.
  • Besides this, the Lean type theory is known not to terminate. Coquand and Abel (Abel and Coquand, 2020) constructed a counterexample to strong normalization using reduction of proofs, and this can be shown to impact definitional equality checks even for regular types:
@@
    @cast (True True’) A
    (propext fun _ => a, fun _ => id⟩)
    (fun z => z (True True’) id z)
@@
    example : foo.recOn (fun _ => 1) = 1 := by
    rfl isDefEq nontermination
@@
  • The kernel does not loop forever in many cases because this is a bad user experience — it has timeouts and depth limits. Not all parts of the kernel have such limits, but it does give us a reasonable design principle which fortuitously solves our termination problem. [^8] There are alternative methods for representing partial functions such as the Bove-Capretta method (take a proof of termination as argument), but we do not want to use this as the kernel API is fixed to match Lean’s actual kernel function, and we also actually want it to time out on extreme cases.
@@ 4. The global environment
  • add : Environment ConstantInfo Environment
    – adds a ConstantInfo declaration to the environment
  • mkEmptyEnvironment : IO Environment
    – constructs an environment with no constants [^9] This is in IO because environment extensions can perform computation in IO for their initialization. But the constant map, which is what we care to verify, is initialized to the empty map.
  • find? : Environment Name Option ConstantInfo
    – retrieves a declaration by name
@@
structure VConstant where
(uvars : Nat) (type : VExpr)
structure VDefEq where
(uvars : Nat) (lhs rhs type : VExpr)
@@
    checks that $e:\alpha$ and then leaves the environment as is.
@@
5. Inductive types
Lean’s implementation of inductive types is based on Dybjer (Dybjer, 1994). It notably differs from Coq in that rather than having primitive fix and match expressions, we have constants called “recursors” generated from the inductive specification. This choice significantly simplifies the VExpr type, which as we have seen contains no special support for anything inductive-related except for the generic def.eq. hook t-extra. It also means we do not need a complex guard checker, something which MetaCoq currently axiomatizes (Sozeau et al., 2019) and which has been the source of soundness bugs in the past. Instead, the complexity is pushed to the generation of the recursor for the inductive type. [^10] It would be premature to draw conclusions about the relative complexity of these two approaches though, as many of the important theorems are unfinished.
For single inductives, the algorithm is described in full in (Carneiro, 2019). In Lean 3, nested and mutual inductives were simulated using single inductives so the kernel only had to deal with the single inductive case. However this simulation process was both costly and did not cover the full grammar of nested and mutual inductives, so Lean 4 moved the checking to the kernel.
For a simple example of a mutual inductive, we can define Even and Odd by mutual induction like so:
@@ 5. Inductive types
| zero : Even 0
| succ : Odd n Even (n+1)
@@
@Even.rec m1 m2 Fz Fs Fs 0 Even.zero Fz
@Even.rec m1 m2 Fz Fs Fs (n+1) (Even.succ e)
@Fs n e (@Odd.rec m1 m2 Fz Fs Fs n e)
@Odd.rec m1 m2 Fz Fs Fs (n+1) (Odd.succ e)
@Fs n e (@Even.rec m1 m2 Fz Fs Fs n e)
@@
{m1 : T Sort u}
{m2 : List T Sort u}
((a : List T) m2 a m1 (T.mk a))
m2 []
@@
m1 head m2 tail m2 (head :: tail))
(t : T) m1 t
@@
Currently, Lean4Lean contains a complete implementation of inductive types, including nested and mutual inductives and all the typechecking effects of this, but not much work has been done on the theoretical side, defining what an inductive specification should generate. Ideally we would like to do it in a way which is not so intimately tailored to the nested inductive collection process (i.e. we can see the order in which subterms are visited by the order of motive_\$n\$ arguments in the above recursor) and instead allows a range of implementation strategies which can all be proved sound against a set-theoretic model.
5.1. Eta for structures
Another new kernel feature in Lean 4 is primitive projections, and eta for structures, which says that if $s$ is a structure (which is to say, a one-constructor non-recursive inductive type), then $s\equiv\mbox{mk}(s.1,\dots,s.n)$ where $s.i$ is the $i$’th projection (Expr.proj). This was already true as a propositional equality proven by induction on $s$, but it is now a definitional equality and the kernel has to unfold structures when needed to support this. This does not fundamentally change the theory; indeed in (Carneiro, 2019) a mild form of eta for pairs was introduced in order to make the reduction to W-types work, so we believe that the main results should still hold even with this extension.
6. Results